The perimeter of a triangle is 153 inches. The length of the longest side is 5/4 the length of the shortest side and the length of the other side is 10 inches more than the length of the shortest side. Find the length of each side.
Answer by scott8148(6628)

let x=(shortest side); so 5x/4=(longest side), and x+10=(other side)

p=x+(5x/4)+(x+10)